OpenStudy (anonymous):

use log to salve exponential equations 0.10*25^x=12.5

OpenStudy (anonymous):

.10*25^x=12.5 25^x=12.5/.010=125 now use log log 25^x=log125 xlog25=log125 x=log125/log25 use calculator lol

OpenStudy (anonymous):

1.5 ans

OpenStudy (anonymous):

is this the answer you need to find out cody?

OpenStudy (anonymous):

good luck now

OpenStudy (anonymous):

thanks mark i get the concept.